6. The power rating of a resistor is the maximum power it can safely dissipate without damage from overheating. A 9.0 k2 resistor is connected across a 120 V potential difference. The power rating for the resistor must be at least 01.6 W 01.1 W 02.8 w 07.5 W
Answer #1

Power P = V^2/R = 120×120/9000

Or P(min)= 1.6W

Here power rating must be atleast 1.6W . It may be more that it but in question least is asked